### v2

@@ 1 Introduction
Emulsions are dispersions of droplets in a suspending continuous liquid phase [1, 2, 3, 4]. Examples of emulsions include food materials like milk, creams, salad dressings, and mayonnaise and cosmetics marketed as lotions and creams. Pharmaceutical formulations like certain eye drops, skin care lotions and oral emulsions are designed such that oil phase serves as a carrier for certain hydrophobic bioactives. Mixing of crude oil and water during petroleum extraction or in oceans after oil spills produces petroleum emulsions. Blends of immiscible polymer solutions or melts that form dispersions of droplets in a suspending liquid are also emulsions. Formulating emulsions with flow properties suitable for processing, applications, and sensory perception involves quests that belong to the realm of rheology, i.e., the science of deformation and flow of simple and complex fluids (or soft matter) [3, 4]. Characterization of emulsion rheology involves the measurement of response to applied stress, strain, or strain rate, typically using specialized equipment called rheometers that are designed to create viscometric flows, or well-defined flow fields [3, 4, 5]. The deformability of drops, the possibility of flow within them, and their coalescence or breakup contribute to flow properties that can be quite distinct from the complex fluids containing dispersed particles, micelles, or macromolecules [1, 2, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Additionally, an emulsion’s stability and flow properties depend on the composition, structure, and mechanical properties of the interface between the dispersed and continuous phases [15, 16, 17, 18, 19]. In this contribution, we highlight how size, shape, concentration, interactions, and interfacial properties of dispersed drops influence droplet concentration-dependent variation in the rheological response of emulsions.
@@
Processing operations like pumping, dispensing, pouring, spreading, and even emulsion stability or shelf-life are influenced by shear viscosity, which characterizes the resistance to shear flows associated with velocity gradients perpendicular to the flow direction [5, 4]. Most published emulsion rheology studies focus on the response to shear flows that commonly arise near solid-liquid interfaces, including flows through channels and around immersed objects and primarily describe magnitude and measurement of viscosity, $\eta$ that quantifies viscous resistance to flow [4, 11, 12, 13, 20, 21, 18]. Stream-wise velocity gradients associated with extensional flows commonly arise in converging channels, porous media, and free-surface flows involving the formation of liquid necks that undergo capillarity-driven pinching [5]. Extensional rheology response profoundly influences processing, applications, and consumer use and perception of emulsions. However, due to longstanding experimental and modeling challenges, studies of extensional rheology of emulsions are less common [22, 23, 24, 25]. Spherical drops deform into ellipsoidal shapes in response to weak velocity gradients [6, 7, 26, 8, 9, 18] and can undergo large deformations in response to strong flows, forming slender bodies and even undergoing capillarity-driven breakup [27, 28, 29, 30]. Emulsification or emulsion formation, mixing and blending liquids, and emulsion rheology are three important class of problems involving drop deformations in response to flow fields [30, 18, 31, 32, 33]. Analytical approaches capture minor or small deformations from spherical shape, but numerical approaches are necessary to model large deformations and breakup or coalescence of drops.
@@
Emulsion drops deformed by velocity gradients display elasticity due to interfacial tension, and after flow stops, drops can recover their unperturbed spherical shape, as it is the minimum energy configuration for fixed drop volume [7]. The characteristic timescale for recovering this interfacial energy-favored state is called relaxation time [7]. The surface tension relaxation time, as it is called sometimes in the emulsion rheology literature, appears in the studies of pinching, coalescence and spreading as viscocapillary time as it captures the interplay of viscous and interfacial stresses [34]. Somewhat analogous elastic response is displayed on stopping of flow by perturbed polymer chains in dilute polymer solution, with a relaxation time providing a measure for the characteristic time over which the entropically favored unperturbed coiled state is recovered [35]. In both dilute emulsions and polymer solutions, this elastic recovery of the unperturbed drop shape or coil configuration is at the heart of viscoelastic behavior, captured as modulus in stress relaxation and oscillatory shear measurements or manifested in rod climbing or steady shear torsional rheometry as non-zero normal stress differences.[4, 5]
@@
In nondilute emulsions and particle suspensions, pairwise and higher order interactions and the local arrangement of discrete drops or particles, referred to as microstructure influence the flow behavior.[4, 5, 11] In-situ visualization or monitoring of microstructural evolution in flow fields by optical or spectroscopic methods shows that the rearrangement of drops and the magnitude of drop deformation and orientation together determine rheological response, including rate variation of shear viscosity and normal stress differences and amplitude and frequency dependent moduli measured using oscillatory shear [18, 30, 27, 28, 36]. In the jammed dense emulsion, the flow behavior is additionally influenced by deformation and flows in interconnected liquid films, leading to a yield stress that must be exceeded before flow can be observed, and typically, viscosity exhibits a deformation rate-dependent or stress-dependent nonlinear response [37, 38, 39, 40].
@@
In this brief review, we highlight theoretical and numerical advances in modeling flows of dilute to dense jammed emulsions. The review is divided into six sections. Section 2 and 3 provide motivation, scope, brief history, definitions, transport equations and dimensional analysis. Section 4 presents small deformation theory and constitutive models for dilute emulsions. Section 5 describes constitutive models and numerical methods developed for non-dilute emulsions. Section 6 is a short survey of jammed dense emulsions, and Section 7 lists a few challenges and opportunities.
@@ 2 Classifying emulsions and mapping concentration-dependent rheology
Classifying emulsions Emulsions are classified using many criteria, ranging from the choice of dispersed and suspending liquid, interface composition, application (food, pharmaceutical, personal care and cosmetics, petroleum) and drop size and volume fraction range [1, 2, 3, 41, 42]. Emulsions are often described based on the choice of dispersed and suspending phase, oil-water or water-oil emulsions that can be obtained by mechanical mixing, phase separation, microfluidics, vapor condensation, or biologically, as in milk. Here, oil can refer to vegetable oils, crude oil (or derived oil), silicone oils, polymerizable monomers (in latex), or even organic liquids, while the water phase can be made with an aqueous solution or water-based mixed solvent. Both milk and mayo are examples of oil-water emulsions, containing water as the suspending or continuous liquid. Distinct from such emulsions are water-in-water emulsions spontaneously formed as complex coacervate forms between two oppositely charged polyelectrolytes and phase separates forming emulsions that are unstable and have short shelf-life [43], though recent studies describe attempts to enhance stability against coalescence [44].
Typical household emulsions like milk, mayonnaise, cosmetic lotions and creams, salad dressings, and fabric softeners appear milky due to scattering by drops with sizes greater than the wavelength of visible light (drop sizes, a > 1 micron). These are examples of macroemulsions and, being thermodynamically unstable, have a finite shelf-life that can be enhanced by reducing drop sizes and size dispersity, diminishing density difference, increasing suspending fluid viscosity and manipulating drop-drop interactions.[1, 2, 3, 14] Like macroemulsions, nanoemulsions (sometimes called miniemulsions) are also thermodynamically unstable, but smaller drop sizes ($a$ = 50-500 nm) and tighter size control lead to prolonged kinetic stability [1, 45, 46, 47]. In contrast, microemulsions that have relatively small drop sizes ($a$ = 10-100 nm) are thermodynamically stable and appear transparent. Classification based on interface composition: surfactant, protein, lipid, particles, polymers, or complexes between these, emphasizes the critical role played by adsorbed species and interfacial rheology on influencing flow properties and stability of emulsions [1, 3, 18].
@@
Concentration-dependent regimes: dilute to dense Constitutive equations that model flow properties of emulsions consider the influence of number density, interactions, and deformation of drops [1, 4, 11, 13, 12, 20, 42]. The exhibited rheological behavior is considered as linear response if the measured flow properties (stress, viscosity or modulus) do not depend on the impelling quantities (stress, strain, or strain rate). Dilute solutions exhibit viscosity or resistance to flow that is comparable to suspending fluid, as can be observed for animal milks, which are examples of emulsions with a relatively low $\phi$. In the dilute regime, the macroscopic properties that capture linear viscoelastic response, including $\eta_{0}$ increase linearly with $\phi$. The deformation and hydrodynamics of each drop in dilute emulsion can be considered independently, by neglecting the influence of hydrodynamic and thermodynamic interactions. In semi-dilute solutions, pairwise interactions make relative viscosity exhibit a nonlinear increase with $\phi$. In concentrated emulsions, drops are so closely packed that drop mobility and deformation become highly restricted by caging or surrounding drops. The shear viscosity exhibits non-Newtonian response for non-dilute emulsions, and elastic effects become progressively stronger with increase in $\phi$. The semi-dilute to highly concentrated emulsions contain a progressively higher $\phi$ (or number density of drops), and influence of associative and repulsive inter-drop interactions and microstructure become manifest and measurable [11, 1, 10, 48].
@@
Figure 1 illustrates that four concentration regimes, dilute, semi-dilute, concentrated and highly concentrated emulsions, can be identified by examining the variation in relative viscosity, $\eta_{r}$ on increasing droplet volume fraction, $\phi$. Here $\eta_{r}=\eta/\mu$ representing the emulsion’s zero shear viscosity scaled with suspending fluid viscosity, $\mu$. Viscosity increases with $\phi$ substantially in the highly concentrated regime, qualitatively emulating the behavior of rigid particle suspensions, where viscosity diverges close to maximum volume fraction [49, 50, 14]. Due to deformability of drops, droplet volume fraction can be increased further leading to the jammed dense emulsion regime. As volume fraction of drops lies beyond the maximum packing fraction for spherical or ellipsoidal particles, jammed dense emulsions contain polygonal-shaped drops separated by interconnected liquid films with a foam-like microstructure. Mayonnaise, an egg-based emulsion of vegetable oil droplets suspended in a aqueous medium [25], is an example of jammed dense emulsion containing closely-packed, polygonal drops, with volume fraction of the drop phase between $\sim 65\%-80\%$. Such dense emulsions display yield stress, $\tau_{Y}$, and elastic modulus, $G$, that increases with volume fraction, [1, 4, 11, 20].The variation in yield stress and modulus scaled by capillary pressure is illustrated in the Figure 1 for jammed dense emulsions. Though it is well-established that increasing drop volume fraction leads to a transition from suspension-like to foam-like behavior as shown schematically, for emulsion drops and for deformable particles, the transition region is dependent on many factors including size and shape, size dispersity, interactions, and mechanisms underlying the deformability of the dispersed phase.[11, 51, 52, 1, 14]
@@
Highlights from 90 years of analytical models for emulsion rheology The review encompasses nearly a century of models that rely on small deformation theory, a perturbation calculation for weak deviations about a spherical shape that are apt for dilute solutions [6, 53, 26, 8, 9, 54, 55]. Taylor (1932) first analyzed drop deformation in the presence of flow [6] and described the viscosity of dilute emulsions by generalizing Einstein’s theory (1906) for a suspension of hard spheres.[56, 57] Decades later, Schowalter, Chaffey and Brenner (1968) [26] extended the model to suggest the existence of normal stress components, but their model reveals no viscosity variation due to drop deformation. Frankel and Acrivos (1970) [9], and Barthès-Biesel and Acrivos (1972) [58] developed constitutive equations for dilute emulsions that describes the response to transient flows. Choi and Schowalter (1975) [10] carried out the extension to semi-dilute solutions, whereas Princen and Kiss (1980s) [37] showed the connection between yield stress or elastic modulus and surface tension for dense emulsions and foams. Flumerfelt (1980) first examined the influence of interfacial tension variation as well as dilatational and shear interfacial viscosity on drop deformation in small deformation limit, and later Leal, Stone and coworkers carried out more extensive examination in the limit large deformation, including the influence of surfactants [19, 59, 54, 29, 18, 16]. Barthès-Biesel (1980) began the examination of deformation and rheology of capsules, defined as viscous drops covered with elastic membranes, and showed that the combination of liquid-like interior enclosed within a solid-like shell leads to behaviors that cannot be inferred from suspension of hard spheres or emulsion containing drops with Newtonian interfaces [60, 61, 62, 63]. Oldroyd (1954, 1955) [7, 64] presented the first attempt at describing the rheology of nondilute emulsions by adopting the effective medium theory proposed in 1946 for a dispersion of deformable particles [65]. Oldroyd also introduced a tensorial framework to capture the complex viscoelastic response of emulsions with appropriate attention to frame invariance. Starting with Taylor’s discussion of drop deformation [6] or with Oldroyd’s framework [7, 64], a large number of analytical and continuum models have been emerged, which incorporate the interplay of drop deformation, interactions, breakup and coalescence processes and rely on numerical and computation approaches especially for connecting microstructure and rheology of nondilute and dense emulsions. We provide selective (and incomplete) but pragmatic overview of the theoretical framework necessary for modeling emulsion rheology.
@@
Scope of this review We provide a brief synopsis of the small-deformation theories based on perturbation methods that are used for capturing drop deformation and rheological response of dilute emulsions to viscometric flows. As dilute emulsions contain non-interacting drops, their shear rheology response under weak flows, including shear viscosity, can be computed by recognizing that contributions from each drop, mildly perturbed drop by the imposed flow must be added to those by the suspending fluid [66, 26, 9, 10, 67, 68, 69, 70, 71, 72, 73, 54, 74, 55]. Deviation from small-deformation conditions are captured by numerical simulations. We organize the discussion according to the composition of the droplet interface in clean drops, and surfactant-covered drops modelled as droplets with surface viscosity. Quantitative descriptions of the rheological response for non-dilute emulsions relies on supplementing analytical theories with computational fluid dynamics to determine the contributions from deformation and dynamics of non-interacting or interacting drops and molecular aspects that control the interfacial properties. We tabulate different methods and highlight their key findings. As the macroscopic rheology response of emulsions is often compared with the expectations of constitutive models developed for suspensions of undeformable particles, we include suitable references for completeness [56, 57].
@@
In this opinion, we exclude discussions relevant to emulsification and highly nonlinear flows of emulsions[31, 32, 33]. We also exclude the discussion of emulsions containing viscoelastic interfaces, drops or suspending liquids [75, 76, 77, 73, 18, 1] and we exclude studies on capsule suspensions [62, 63]. We cite a paucity of datasets and the immensity of challenges involved in theoretical and experimental studies of extensional rheology response as a reason for excluding the few published studies, including our own [25, 23, 78, 79]. We do not cover studies on Pickering emulsions, water-in-water emulsions, microemulsions, and nanoemulsions, and recommend some recent reviews [1, 45, 46, 80, 44, 81]. We exclude any discussion of rheometry techniques and measured rheological response of emulsions or of interfaces enriched with adsorbed species but we anticipate the references included can be used as a guidebook for the road not taken [1, 2, 4, 82, 11, 12, 5, 83, 17, 84, 85]. Capillary pressure, interfacial rheology, disjoining pressure (contributed by intermolecular and surface forces), and bulk rheology of two liquids all influence drainage flows in liquid films separating any pair of deformed drops, and though interplay and drainage kinetics affect emulsions stability and rheology, a comprehensive description of these remains an open challenge [1, 11, 48, 86]. Nevertheless, we plan to highlight reviews, monographs, articles, and textbooks that form essential reading for appreciating state-of-the-art understanding and progress in the experimental, theoretical, and computational studies of emulsion rheology [4, 11, 12, 21, 20, 13, 30, 18, 3].
3 Microhydrodynamics of emulsions: the governing equations and dimensional analysis

### v3

@@ 1 Introduction
Emulsions are dispersions of droplets in a continuous suspending liquid phase [1, 2, 3, 4]. Examples of emulsions include food materials such as milk, creams, salad dressings, chocolate, and mayonnaise and cosmetics marketed as lotions and creams. Pharmaceutical formulations like certain eye drops, skin care lotions, and oral emulsions are designed such that oil phase serves as a carrier for certain hydrophobic bioactives. Dispersion of crude oil in water or vice versa during petroleum extraction or in oceans after oil spills produces petroleum emulsions. Blends of immiscible polymer solutions or melts that form dispersions of droplets in a suspending liquid are also emulsions. The immiscibility implies that the free energy of the mixing is higher than the phase separated systems formed by drops in a matrix phase [4]. The formulation of emulsions with flow properties suitable for processing, applications, and sensory perception involves quests that belong to the realm of rheology, i.e., the science of deformation and flow of simple and complex fluids (or soft matter) [3, 4, 5, 6]. Emulsion rheology is characterized by measuring the response to applied stress, strain, or strain rate, typically using specialized equipment called rheometers that are designed to create viscometric flows, or well-defined flow fields to assess flow behavior [3, 4, 5]. The deformability of drops, the possibility of flow within them, and their coalescence or breakup contribute to emulsion rheology [1, 2, 3, 4] that can be quite distinct from the rheology of complex fluids containing dispersed particles, micelles, or macromolecules [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Furthermore, the stability and flow behavior of emulsions depend on the composition, structure, and mechanical properties of the interface between the dispersed and continuous phases [17, 18, 19, 20, 21]. In this contribution, we highlight how size, shape, concentration, interactions, and interfacial properties of dispersed drops influence droplet concentration-dependent variation in the rheological response of emulsions.
@@
Processing operations such as pumping, dispensing, pouring, spreading, and even emulsion stability or shelf-life are influenced by shear viscosity $\eta$, which characterizes resistance to shear flows associated with velocity gradients perpendicular to the flow direction [5, 4]. Such shear flows commonly arise near solid-liquid interfaces, including pressure-driven flows through channels, and drag flows around immersed objects or near moving substrates. Most published emulsion rheology studies primarily describe the magnitude and measurement of shear viscosity, $\eta$, with a focus on stress or shear rate dependent variation [4, 8, 22, 10, 23, 24, 20]. However, stream-wise velocity gradients associated with extensional or elongational flows commonly arise in converging channels, porous media, and free surface flows, involving the formation of liquid necks that undergo capillarity-driven pinching [5, 4]. Studies of extensional rheology of emulsions are less common due to longstanding experimental and modeling challenges [25, 26, 27, 28]. Extensional rheology responses profoundly influence the processing, applications, and consumer use and perception of emulsions, described in heuristic terms such as sprayability, jettability, stringiness, ropiness, and jettability [25, 26, 27, 28].
Spherical drops deform into ellipsoidal shapes in response to weak velocity gradients [11, 12, 13, 14, 15, 20] and can undergo large deformations in response to strong flows, forming slender bodies and even undergoing capillarity-driven pinching leading to breakup [29, 30, 31, 32]. Emulsification or emulsion formation, liquid blending, and emulsion rheology are three important problems that involve drop deformations in response to flow fields [32, 20, 33, 34, 35, 36]. Analytical approaches capture minor or small deformations from spherical shape, but numerical approaches are necessary to model large deformations and pinching, breakup or coalescence of drops, especially, for emulsions containing the dispersed drop phase in higher volume fractions.
@@
Emulsion drops deformed by velocity gradients display elasticity due to restoring stresses set by interfacial tension. After flow stops, drops can recover their unperturbed spherical shape, as it is the minimum energy configuration for a fixed drop volume [12]. The characteristic timescale for recovering this interfacial energy-favored state is called relaxation time [12] or shape or surface tension relaxation time in the emulsion rheology literature [9, 4, 20]. The shape relaxation time appears as viscocapillary time in interfacial fluid mechanics, including the studies of pinching, coalescence, and spreading of drops [37, 38] as it captures the interplay of viscous and interfacial stresses. Somewhat analogous elastic response is displayed in dilute polymer solutions by polymer chains perturbed by flow, with a relaxation time as the characteristic time over which the unperturbed, entropically favored coiled state is recovered after the cessation of the flow [39, 4]. In both dilute emulsions and polymer solutions, this elastic recovery of the unperturbed drop shape or coil configuration is at the heart of viscoelastic behavior, captured as modulus in stress relaxation and oscillatory shear measurements, or manifested in rod climbing or steady shear torsional rheometry as stresses due to nonzero normal stress differences [4, 5].
@@
In non-dilute emulsions and particle suspensions, pairwise and higher order interactions and the local arrangement of discrete drops or particles constitute the microstructure that influences the flow behavior [4, 5, 6, 8]. In-situ visualization or monitoring of the evolution of microstructure in flow fields by optical or spectroscopic methods shows that the rearrangement of drops and the magnitude of drop deformation and orientation together determine the rheological response, including rate variation of shear viscosity and normal stress differences and amplitude and frequency dependent moduli measured using oscillatory shear [20, 32, 29, 30, 40]. In the jammed dense emulsion, the shear flow behavior is also influenced by deformation and flows in interconnected liquid films, leading to a yield stress that must be exceeded before flow can be observed, and typically, shear viscosity exhibits a deformation rate- or stress- dependent nonlinear response [41, 42, 43, 44]. Due to the enhanced nonlinearity and complexity of the problems, few studies explore the response of the non-dilute emulsions in extensional flows and confined flows [45]. Far fewer theoretical and simulation studies account for the influence of non-Newtonian response (rate-dependent shear and extensional viscosity, transient and nonlinear viscoelasticity) of the suspending or dispersed liquid or of the decorated, populated interface [45].
@@
In this brief review, we highlight theoretical and numerical advances in modeling flows of dilute to dense jammed emulsions. The review is divided into six sections. Section 2 and 3 provide motivation, scope, brief history, definitions, transport equations, and dimensional analysis. Section 4 presents the small deformation theory and constitutive models for dilute emulsions. Section 5 describes constitutive models and numerical methods developed for non-dilute emulsions. Section 6 is a short survey of jammed dense emulsions, and Section 7 lists a few challenges and opportunities. We have included a primer on the small deformation theory in Appendices A and B, retracing steps taken by Frankel and Acrivos [15] from the shape distortion tensor to the constitutive equation for dilute emulsions that can describe both transient and steady state non-Newtonian and viscoelastic response.
@@ 2 Classifying emulsions and mapping concentration-dependent rheology
Classifying emulsions Emulsions are classified using many criteria, ranging from the choice of dispersed and suspending liquid, interface composition, application (food, pharmaceutical, personal care and cosmetics, petroleum) and drop size and volume fraction range [1, 2, 3, 46, 47]. Emulsions are often described on the basis of the choice of dispersed and suspending phase, oil-water or water-oil emulsions that can be obtained by mechanical mixing, phase separation, microfluidics, vapor condensation, or biologically, as in milk. Here, oil can refer to vegetable oils, crude oil (or derived oil), silicone oils, polymerizable monomers (in latex), or even organic liquids, while the water phase can be made with an aqueous solution or water-based mixed solvent. Both milk and mayo are examples of oil-water emulsions, containing water as the suspending or continuous liquid. Unlike such emulsions, water-in-water emulsions spontaneously appear as complex coacervate forms between two oppositely charged polyelectrolytes and phase separates forming emulsions that are unstable and have short shelf-life [48], though recent studies describe attempts to enhance stability against coalescence [49].
Typical household emulsions, such as milk, mayonnaise, cosmetic lotions and creams, salad dressings, and fabric softeners appear milky due to scattering by drops with sizes greater than the wavelength of visible light (drop sizes, a > 1 micron). These are examples of macroemulsions and, being thermodynamically unstable, have a finite shelf-life that can be enhanced by reducing drop sizes and size dispersity, diminishing density difference, increasing the suspending fluid viscosity and manipulating drop-drop interactions [1, 2, 3, 7]. Like macroemulsions, nanoemulsions (sometimes called miniemulsions) are also thermodynamically unstable, but smaller drop sizes ($a$ = 50-500 nm) and tighter size control lead to prolonged kinetic stability [1, 50, 51, 52]. In contrast, microemulsions that have relatively small drop sizes ($a$ = 10-100 nm) are thermodynamically stable and appear transparent. Classification based on interface composition: surfactant, protein, lipid, particles, polymers, or their complexes emphasizes the critical role that adsorbed species and interfacial rheology play in influencing the flow properties and stability of emulsions [1, 3, 20].
@@
Concentration-dependent regimes: dilute to jammed dense Constitutive equations that model flow properties of emulsions consider the influence of number density, interactions, and deformation of drops [1, 4, 8, 10, 22, 23, 47]. The exhibited rheological behavior is considered as a linear response if the measured flow properties (stress, viscosity, or modulus) do not depend on the impelling quantities (stress, strain, or strain rate). Dilute emulsions exhibit viscosity or resistance to flow that is comparable to suspending fluid, as can be observed for animal milks, which are examples of emulsions with a relatively low $\phi$. In the dilute regime, the macroscopic properties that capture the linear viscoelastic response, including $\eta_{0}$ increase linearly with $\phi$. The deformation and hydrodynamics of each drop in dilute emulsion can be considered independently, by neglecting the influence of hydrodynamic and thermodynamic interactions. In semi-dilute emulsions, pairwise interactions make relative viscosity exhibit a nonlinear increase with $\phi$. In concentrated emulsions, drops are so closely packed that drop mobility and deformation become highly restricted by caging or surrounding drops. The shear viscosity exhibits a stronger non-Newtonian response for the non-dilute emulsions, and the elastic effects become progressively stronger with an increase in $\phi$. The semi-dilute to highly concentrated emulsions contain a progressively higher $\phi$ (or number density of drops), and influence of associative and repulsive inter-drop interactions and microstructure become manifest and measurable [8, 1, 16, 53].
@@
Figure 1 illustrates that four concentration regimes, dilute, semi-dilute, concentrated and highly concentrated emulsions, can be identified by examining the variation in relative viscosity, $\eta_{r}$ on increasing droplet volume fraction, $\phi$. Here $\eta_{r}=\eta/\mu$ representing the emulsion’s zero shear viscosity scaled with suspending fluid viscosity, $\mu$. Viscosity increases with $\phi$ substantially in the highly concentrated regime, qualitatively emulating the behavior of rigid particle suspensions, where viscosity diverges close to maximum volume fraction [54, 55, 7]. Due to deformability of drops, droplet volume fraction can be increased further leading to the jammed dense emulsion regime. As volume fraction of drops lies beyond the maximum packing fraction for spherical or ellipsoidal particles, jammed dense emulsions contain polygonal-shaped drops separated by interconnected liquid films with a foam-like microstructure. Mayonnaise, an egg-based emulsion of vegetable oil droplets suspended in a aqueous medium [28], is an example of jammed dense emulsion containing closely-packed, polygonal drops, with volume fraction of the drop phase between $\sim 65\%-80\%$. Such dense emulsions display yield stress, $\tau_{Y}$, and elastic modulus, $G$, that increases with volume fraction [1, 4, 8, 23].The variation in yield stress and modulus scaled by capillary pressure is illustrated in the Figure 1 for jammed dense emulsions. Though it is well-established that increasing drop volume fraction leads to a transition from suspension-like to foam-like behavior as shown schematically, for emulsion drops and for deformable particles, the transition region depends on many factors including size and shape, size dispersity, interactions, and mechanisms underlying the deformability of the dispersed phase [8, 56, 57, 1, 7].
@@
Highlights from ninety years of analytical models for emulsion rheology The review encompasses and primarily highlights models that rely on small deformation theory, a perturbation calculation for weak deviations about a spherical shape that are apt for dilute emulsions, and provide insights into rheology of nondilute emulsions [11, 58, 14, 15, 59, 60, 9, 12, 10]. In 1906, Einstein connected the microhydrodynamics calculation of the flow around dispersed particles to the estimation of viscosity of a dilute suspension of hard spheres [61, 62]. Subsequently, Taylor (1932) first analyzed drop deformation in the presence of flow [11] and generalized Einstein’s theory [61, 62] to describe the viscosity of dilute emulsions by accounting for internal circulation. Decades later, Schowalter, Chaffey and Brenner (1968) [13] extended the model to suggest the existence of normal stress components, but their model reveals no viscosity variation due to drop deformation. Frankel and Acrivos (1970) [15], and Barthès-Biesel and Acrivos (1972) [63] developed constitutive equations for dilute emulsions that describe the response to transient flows. Choi and Schowalter (1975) [16] carried out the extension to semi-dilute solutions, whereas Princen and Kiss (1980s) [41] showed the connection between yield stress or elastic modulus and surface tension for dense emulsions and foams. Flumerfelt (1980) first examined the influence of interfacial tension variation as well as dilatational and shear interfacial viscosity on drop deformation in the small deformation limit [21]. Later, Leal, Stone and coworkers performed more extensive examination in the limit of large deformation, including the influence of surfactants [64, 59, 31, 20, 18].
Barthès-Biesel (1980) began the examination of deformation and rheology of capsules, defined as viscous drops covered with elastic membranes, and showed that the combination of liquid-like interior enclosed within a solid-like shell leads to behaviors that cannot be inferred from suspension of hard spheres or emulsion containing drops with Newtonian interfaces [65, 66, 67, 68]. Oldroyd (1954, 1955) [12, 69] presented the first attempt to describe the rheology of nondilute emulsions by adopting the effective medium theory proposed in 1946 by Fröhlich and Sack for the dispersion of deformable particles [70]. Oldroyd also introduced a tensorial framework to capture the complex viscoelastic response of emulsions with appropriate attention to frame invariance. Starting with Taylor’s discussion of drop deformation [11] or with Oldroyd’s framework [12, 69], a large number of analytical and continuum models have emerged, which incorporate the interplay of drop deformation, interactions, breakup and coalescence processes and rely on numerical and computation approaches, especially for connecting the microstructure and rheology of nondilute and dense emulsions. We provide a selective (and incomplete) but pragmatic overview of the theoretical framework necessary for modeling emulsion rheology.
@@
Scope of this review We provide a brief synopsis of the small-deformation theories based on perturbation methods that are used to capture drop deformation and the rheological response of dilute emulsions to viscometric flows. As dilute emulsions contain non-interacting drops, their shear rheology response under weak flows, including shear viscosity, can be computed by recognizing that contributions from each drop, the mildly perturbed drop by the imposed flow, must be added to those by the suspending fluid [71, 13, 15, 16, 72, 73, 74, 75, 76, 77, 9, 59, 78, 60, 14]. Deviations from small-deformation conditions are captured by numerical simulations. We organized the discussion according to the composition of the droplet interface in clean drops and surfactant-covered drops modeled as droplets with surface viscosity. Quantitative descriptions of the rheological response for non-dilute emulsions rely on supplementing analytical theories with computational fluid dynamics to determine the contributions from deformation and dynamics of non-interacting or interacting drops and molecular aspects that control the interfacial properties. We tabulate different methods and highlight their key findings. As the macroscopic rheological response of emulsions is often compared with the expectations based on constitutive models developed for suspensions of undeformable particles, we include suitable references for completeness [61, 62, 4, 54, 6, 7, 5].
@@
In this opinion, we exclude discussions relevant to emulsification and highly nonlinear flows of emulsions [33, 34, 35]. We briefly mention the influence of viscoelastic interfaces, drops, or suspending liquids on the emulsion rheology and for the sake of brevity, highlight articles and reviews that detail recent progress and open questions [79, 80, 45, 9, 20, 1]. Likewise, we exclude studies on capsule suspensions [67, 68], wall effects, and the influence of external force fields on droplet topology and emulsion microstructure [81, 82, 83, 84, 85, 45, 86, 87]. We cite a paucity of datasets and the immensity of challenges involved in theoretical and experimental studies of the extensional rheology response as a reason for excluding a detailed exposition of the few published studies, including our own [28, 26, 88, 89]. We do not cover studies on Pickering emulsions, water-in-water emulsions, microemulsions, and nanoemulsions, and recommend some recent reviews [1, 50, 51, 90, 49, 91]. We exclude any discussion of rheometry techniques and measured rheological response of emulsions or interfaces enriched with adsorbed species, but we anticipate that the references included here can be used as a guide for the road not taken [1, 2, 4, 92, 8, 22, 5, 93, 19, 94, 95, 3, 20, 21, 23]. Although capillary pressure, interfacial rheology, disjoining pressure (contributed by intermolecular and surface forces), and bulk rheology of two liquids all influence drainage flows in thin liquid films that separate any droplet pairs and therefore influence emulsion stability and rheology, a comprehensive description of these remains an open challenge [1, 8, 53, 96]. However, we plan to highlight reviews, monographs, articles, and textbooks that form essential reading for appreciating the state-of-the-art understanding and progress in experimental, theoretical, and computational studies of emulsion rheology [4, 8, 22, 23, 10, 9, 20, 3, 21, 18].
3 Emulsion microhydrodynamics: the governing equations and scaling